The decomposition of N2O5 can be described by the equation

2N2O5(soln)⟶4NO2(soln)+O2(g)

Given the data in the table for the reaction at 45 °C in carbon tetrachloride solution, calculate the average rate of reaction for each successive time interval.

t (s) [N2O5] (M)
0 2.463
165 2.220
506 1.792
815 1.475

interval: 0 s to 165 s average rate of reaction: M/s

interval: 165 s to 506 s average rate of reaction: M/s

interval: 506 s to 815 s average rate of reaction: M/s

Nitrogen pentoxide decomposes to produce nitrogen dioxide and oxygen according to the equation.

2N2O5(g)⟶4NO2(g)+O2(g)2N2O5(g)⟶4NO2(g)+O2(g)

The following data were collected for this reaction at 298K. Calculate the rate of change for N2O5 between 100100 s and 300300 s.

Time (s) [N2O5] (M) [NO2] (M) [O2] (M)
0 0.0200 0 0
100 0.0169 0.0063 0.0016
200 0.0142 0.0115 0.0029
300 0.0120 0.0160 0.0040
500 0.0086 0.0229 0.0057
700 0.0061 0.0278 0.0070

M/s
Answer #1

Answer:

I. The rate can be written as;

Rate of disappearance of N2O5, rd = d[N2O5]/dt

Rate of reaction = -1/2 x rd

For 0 - 165 s

rateavg = - 1/2 x (2.220 - 2.463)/(165 - 0) = 7.36 x 10-4 M/s

For 165 - 506 s

rateavg = - 1/2 x (1.792 - 2.220)/(506 - 165 ) = 6.27 x 10-4 M/s

For 506 - 815 s

rateavg = - 1/2 x (1.475 - 1.792)/(815 - 506) = 5.12 x 10-4 M/s

II. Similar to the above question, the rate of change of N2O5 can be given by the simple relation;

rd = d[N2O5]/dt = (0.0120 - 0.0169)/(300-100) = -2.45 x 10-5 M/s

Please let me know, if you have any doubt by commenting below the answer or needed more help on the same question.

Thanks